# Community-Cross-Instruct: Unsupervised Instruction Generation for Aligning Large Language Models to Online Communities

### v1

@@ 4 CommSurvey: A Collection of Community-Specific Survey Questions
To measure the alignment of the finetuned foundational LLM to the community, we administer a survey of multi-choice questions to the model and evaluate the agreement between the model’s responses and the community members’ responses. However, manually designing surveys and collecting human responses from online communities is a non-trivial process, which is costly and time-consuming. Instead, we curate another dataset called CommSurvey for evaluating the alignment of the finetuned LLM’s views to the community using survey questions. CommSurvey consists of automatically generated multi-choice survey questions and answers, where each question is paired with answers from different communities (Figure 2(b)). CommSurvey-Pol and CommSurvey-Fit are the two subsets for politics and fitness respectively, consisting of 4,457 and 2,930 multi-choice questions respectively, covering a wide range of topics.
politics r/Anarchism
r/Anarcho_
Capitalism
r/Conservative r/Libertarian r/Republican r/democrats
0.75 0.60 0.75 0.45 0.75 0.65
fitness r/Fitness r/loseit r/keto r/xxfitness r/GettingShredded
0.75 0.70 0.60 0.50 0.65
Table 1: Agreement between human annotators’ survey responses and the advanced LLM (Claude-3) generated survey answers in different domains and subreddits, measured by accuracy.
4.1 Multi-choice Survey Generation.
We initialize an empty multi-choice survey pool. Then, following the open-ended instruction-answer pairs generation in §$\S$3.3, we iteratively generate multi-choice questions and answers (Figure 2(b)) using the same advanced LLM (Claude-3) and add them to the pool. We also explored other advanced LLMs, such as GPT-4 or Gemini, but none of them was able to generate high-quality questions or answers as Claude-3. The prompting template is shown in Figure 8. We assume that the answers generated by the advanced LLM are the “semi-ground truths”, and that they faithfully represent the views of the corresponding communities.
Note that there is a correspondence between instructions in CommInst and survey questions CommSurvey, as they are generated using nearly identical prompts. Following the correspondence, we split CommSurvey-Pol and CommSurvey-Fit following the same two ways for CommInst. Due to the correspondence, there is no information leakage between CommInst-Pol-train and CommSurvey-Pol-eval, or between CommInst-Pol-train-topic and CommSurvey-Pol-eval-topic.
4.2 Validating Generated Survey Answers
We verify the faithfulness of the “semi-ground truths” for the communities via human annotation. For each domain, we randomly sample 20 questions from the survey (see Fig. 9 and Fig. 10 in the Appendix) and ask two expert annotators to answer them from the perspective of the subreddit. We also tried directly posting the surveys to the subreddits and having community members answer them. However, our posts were immediately removed by moderators and our accounts were banned from certain subreddits. This further indicates the difficulty of directly surveying online communities.
[…]
@@ 5 Experiments
We demonstrate the effectiveness of the pipeline on the Reddit politics and fitness domains. We align LLMs to Reddit communities by finetuning them on relevant demonstrations from CommInst. After finetuning, we administer surveys from CommSurvey and evaluate the finetuned LLMs by comparing their responses to the “semi-ground truths”, in order to measure alignment.
@@ 5.1 Experimental Setup
We experiment with two foundational LLMs – GPT-3.5 and Llama-3.
@@
We finetune the foundational LLMs on CommInst-Pol-train and CommInst-Fit-train. We use the OpenAI finetuning API to finetune gpt-3.5-turbo-0125 for 1 epoch, as we find that the loss converges within 1 epoch. The batch size is automatically determined by the API, and the finetuning takes ~40mins. For Llama-3, the implementation is based on LlamaFactory Zheng et al. (2024), and we finetune Llama-3-8B-Instruct with batch size 16 for 2 epochs on 4 H100 GPUs. The finetuning takes ~10mins.
@@
We administer survey questions in CommSurvey-Pol-veal and CommSurvey-Fit-veal to the finetuned foundational LLMs. In the prompt, we provide the question and the four options verbatim, and then add the instruction “Select only one answer by stating either A, B, C, or D. Do not provide any additional explanation or rationale for your choice.”, for the sake of easier matching of the model’s response to the option. For GPT-3.5, the temperature is set to 0.9 and the model provides 5 generations for each question. For Llama-3, temperature is set to 0.9, $top\_p$ is set to 0.9, and the model provides 10 generations for each question. The majority vote from the different generations for each question is taken. We compare the model responses to the “semi-ground truths” and report the accuracy as a measure of alignment.
@@ 5.2 Baselines
Vanilla LLMs.
We administer survey questions to the vanilla unfinetuned GPT-3.5 and Llama-3 and evaluate their answers.
Vanilla LLMs+Steering.
When prompting the vanilla language models to answer survey questions, we steer them to mimic the community, by specifying in the prompt that “Select only one answer that best aligns with the opinions of members from subreddit r/[subreddit]”. Steering the vanilla LLMs can nudge them to respond for the community. However, it is important to note that steering only applies to predefined communities that are developed via a predefined tag, such as subreddits, where we can easily specify community names in the prompt to steer the LLMs. Although in this paper we focus on communities associated with online forums, Community-Cross-Instruct generalizes to other online communities, including organically formed communities in the retweet network Chu et al. (2024) or the news co-sharing network Mosleh et al. (2021); He et al. (2024b), as long as their texts are readily available. It is non-trivial to summarize the organically-formed communities concisely in language, but we can still align LLMs to them via Community-Cross-Instruct.
@@ 5.3 Main Results
We compare survey answers generated by the finetuned LLMs to the “semi-ground truth” and report the accuracy in Figure 4 as a measure of alignment with the corresponding community. In most cases, steering LLMs towards specific communities creates better alignment by relying on the models’ pre-existing knowledge about the communities. However, such pre-knowledge may be incomplete, inaccurate, or outdated. Therefore, by finetuning LLMs on instructions distilled from the text corpora of communities, Community-Cross-Instruct outperforms steering and achieves the best results, demonstrating its effectiveness in aligning LLMs to different communities.
(a) politics
(b) fitness.
Figure 4: Alignment of finetuned LLMs to the corresponding communities measured by accuracy, by comparing the LLMs’ survey responses to the “semi-ground truths”.
The base models show worse alignment with r/Conservative and r/Republican than other Reddit communities. This observation conforms to findings by previous works Santurkar et al. (2023); He et al. (2024a); Chen et al. (2024b) that LLMs tend to display a liberal bias and poor alignment with conservatives. After explicitly steering them towards specific communities, their alignment with the two subreddits significantly increases, but it is still lower than that with other communities. Finetuning the LLMs with Community-Cross-Instruct further improves the alignment, although it is still lower than for other communities. This observation highlights the risks of deploying LLMs in political scenarios due to the entrenched political biases.
For the r/Anarchism subreddit, GPT-3.5-CrossInst achieves similar alignment as the steering baseline. We suspect that this is because views about anarchism are captured by the pre-existing knowledge of GPT-3.5. Therefore, infusing such knowledge again by Community-Cross-Instruct barely leads to any gain.
For the r/Fitness and r/xxfitness communities, steering both LLMs worsens the alignment compared to the vanilla model. r/Fitness is a general fitness subreddit and r/xxfitness is a fitness community for women, both of which may have a broader and more diverse range of opinions compared to other subreddits. Steering the models to align with a singular “community opinion” might oversimplify the nuanced perspectives within these fitness communities, leading to decreased alignment accuracy. However, Community-Cross-Instruct effectively distills knowledge from the community corpora during finetuning. On r/GettingShredded, Community-Cross-Instruct is outperformed by the steering baseline for GPT-3.5, but not Llama-3. r/GettingShredded has the least amount of comments and posts, producing much fewer demonstrations for finetuning (Appendix 3). The insufficient data might be less effective for finetuning a larger LLM like GPT-3.5111We assume that GPT-3.5 has more than 8B parameters., leading to inferior performance.
@@ 5.4 Out-of-Topic Generalizability
Our final goal is to create LLMs aligned to different communities, which can be used to answer any survey question from the perspective of the community. In the real world, the surveys may contain questions covering topics that do not appear in the finetuning data CommInst. In this study, we finetune Llama-3 on CommInst-Pol-train-topic, and evaluate it on CommSurvey-Pol-eval-topic, to make sure that there is no overlap in topic coverage. The results in Fig. 5 show that Community-Cross-Inst consistently outperforms both baselines, demonstrating its superiority in aligning with views of different communities, and generalize the learned views to unseen topics.
Figure 5: Cross-topic alignment of finetuned LLMs to the corresponding communities. The LLMs are finetuned on demonstrations and evaluated on survey questions on different topics.
@@ 6 Related Works
Self-Improved Instruction Tuning. A survey of instruction-tuning approaches Zhang et al. (2023) outlines several methods for models to autonomously self-improve their instruction set. These include generating instructions for pre-existing texts Li et al. (2023), eliciting interaction between different model iterations Chen et al. (2024c), and bootstrapping from an existing instruction set to generate new ones Wang et al. (2022). Additionally, a human-in-the-loop framework builds upon self-generalization by iterating between human and machine-generated instructions Guo et al. (2024). Despite these advances, these strategies require manual input. In this work, we build on self-instruct, a method to enhance instruction tuning by eliciting synthetic model instruction generations. We alter the original framework by removing the required set of manually written seed instructions Wang et al. (2022). This work significantly contributes to the field by offering a scalable and fully autonomous solution to instruction tuning, paving the way for more adaptive and intelligent models.
Community Alignment. Aligning with the dialects and behaviors of online communities is crucial for accurately modeling subpopulations Dorn et al. (2023, 2024). Subpopulation representative models (SRMs) Simmons and Hare (2023) can be used to emulate some characteristics of a particular subpopulation, particularly as LLMs can provide fine-grained, demographically-correlated outputs Argyle et al. (2023a). Using self-instruct with questions from the American Trans Panel as the seed instruction set, LLMs can be ideologically manipulated to reflect particular political opinions Chen et al. (2024b). To learn about partisan communities, Jiang et al. (2022) repeatedly prompt GPT 2 with tweets authored by prominent community members to generate synthetic responses to election survey questions. Argyle et al. (2023b) find that exposing GPT 3 to thousands of socio-demographic backstories leads the model to obtain a complex understandings of sociological concepts. ‘Media diet models’ train BERT using MLM on news articles, evaluating the alignment of model predictions with a survey of community members who follow the media Chu et al. (2023). Feldman et al. (2022) expose GPT 3 to Yelp reviews to expand insights past sentiment information. Prompting language models with (survey question, individual ID, survey year) triplets from the Americans in the General Social Surveys has helped social scientists impute missing survey data and predict opinions Kim and Lee (2023). These techniques rely on manually created datasets at least as a seed dataset to align the models with the desired communities, limiting SRBMs to high-resource environments.
Surveys. Social scientists use surveys to systematically collect data from populations to characterize their beliefs, attitudes, opinions, and behaviors. While designing the survey, researchers consider the order, wording, and format of questions (open-ended vs multiple-choice) to increase clarity and ensure the validity and reliability of the survey data. Survey design also aims to minimize biases, such as the non-response bias Hill et al. (1997), fatigue due to overly long surveys Choi and Pak (2005), social desirability bias Gordon (1987) when asking survey participants questions about sensitive topics (due to the desire to conform to perceived social expectations), or those involving social stigmas Goel and Salganik (2010). These issues, as well as the variety of existing biases, call for new approaches to survey design.

### v2

@@ 4.3 Multi-choice Surveys Generation
To measure the alignment of the finetuned foundational LLM to the community, we administer a survey of multi-choice questions to the model and evaluate the agreement between the model’s responses and the community members’ responses. However, manually designing surveys and collecting human responses from online communities is a non-trivial process, which is costly and time-consuming. Instead, we curate another collection of datasets called CommSurvey for evaluating the alignment of the finetuned LLMs’ views to the communities.
We initialize $n$ empty multi-choice question pools $\{\textsc{CommSurvey}_{i}\}_{i=1}^{n}$. Then, following the open-ended instruction-response pairs generation process in §$\S$4.2, we iteratively generate multi-choice questions and answers (Figure 2(b)) using the same advanced LLM (GPT-4o) and add them to the pools [^4] In practice we use the advanced LLM to generate open-ended instructions and multi-choice questions at the same time in a single query. For clarity of presentation, we articulate them separately.. The prompting template is shown in Appendix B.2. Each question pool $\textsc{CommSurvey}_{i}$ is used to evaluate a finetuned foundational LLM on community $C_{i}$. Please refer to Appendix B.3 for more details about CommSurvey.
We assume that the answers generated by the advanced LLM are the “semi-ground truths”, and that they faithfully represent the views of the corresponding communities. As an empirical evaluation, we compute the pairwise agreement between different communities using their responses to questions in CommSurvey. The results are presented in Figure 5 in Appendix B.4. We observe significant polarization between the left-leaning and right-leaning communities, and strong agreement between the two right-leaning communities (r/Conservative and r/AskThe_Donald). However, in the diet domain where the communities are more practical, there is less polarization. For more rigorous evaluation, we verify the “semi-ground truths” by human annotation, as detailed in Appendix C.
@@ 5 Experiments
We align foundational LLMs to Reddit communities by finetuning them on relevant demonstrations from CommInst. After finetuning, we evaluate the models using surveys from CommSurvey, comparing their responses to “semi-ground truths” to assess alignment.
@@ 5.1 Experimental Setup
Data Generation.
We use OpenAI’s GPT-4o Batch API to generate the datasets. For each query, three open-ended instructions and two multi-choice questions are created, as shown in Appendix B.2. The API costs approximately \$2 for the politics domain and \$5 for the diet domain.
Data Splitting.
Among all the queries used to prompt the advanced LLM to generate data, we randomly select 85% of them and use the instructions generated from these queries as the training instructions denoted as CommInst-train. For the rest 15% queries, we use the corresponding multi-choice questions as test questions, denoted as CommSurvey-test. In addition to randomly splitting the queries, we perform a topic-wise split, and make sure that the 85% training instructions, CommInst-train-topic, and 15% test questions, CommSurvey-test-topic, do not cover the same topics.
@@
We focus on two strong foundational LLMs – Llama-3.1-8B-Instruct Dubey et al. (2024) and GPT-3.5-Turbo Ouyang et al. (2022). For each community $C_{i}$, we finetune the LLM on CommInst-traini (§$\S$5.3) and CommInst-train-topici (§$\S$5.4). The input and the output are the instruction and response verbatim. Llama-3.1 is finetuned with LlamaFactory Zheng et al. (2024), using both full and LoRA finetuning Hu et al. (2022), with batch size 16 for 3 epochs. We report the results of the model (full or LoRA) that achieves better loss on the validation set (5% of the training data). Full finetuning takes around 3 minutes on 4 NVIDIA H100 GPUs, and LoRA finetuning takes around 30 seconds on 1 GPU. For GPT-3.5, we use the OpenAI API, which completes finetuning in around 10 minutes for 1 epoch, where the batch size is automatically determined.
@@
We administer survey questions from CommSurvey-test (§$\S$5.3) and CommSurvey-test-topic (§$\S$5.4) to the finetuned foundational LLMs. Each prompt includes the question and options verbatim, followed by the instruction, “Select only one answer by stating either A, B, C, or D. Do not provide any additional explanation or rationale for your choice,” to facilitate easier matching of the model’s responses to the options. We set the temperature to 0.8, generating 20 responses per question, and take the majority vote. Accuracy is calculated by comparing the model’s answers to the “semi-ground truths”.
Subreddit Llama-3.1-8B GPT-3.5-Turbo
Without Steering With Steering Without Steering With Steering
Context CrossInst Steering Steering+CrossInst Context CrossInst Steering Steering+CrossInst
r/Liberal 8.3 54.2 55.8 58.3 41.7 62.5 45.8 62.5
r/NeutralPol 3.8 50.0 55.0 63.3 55.0 55.0 40.0 50.0
r/Anar_Cap 54.1 76.7 70.0 66.7 50.0 66.7 66.7 73.3
r/Conservative 18.4 63.3 60.5 56.7 50.0 53.3 53.3 53.3
r/AT_Donald 23.7 56.7 66.7 70.0 30.0 56.6 50.0 63.3
avg. politics 21.7 60.2 61.6 63.0 45.3 58.8 51.2 60.5
r/keto 26.0 75.0 67.0 72.0 56.0 68.0 66.0 66.0
r/WLAdvice 27.7 60.6 65.2 61.7 59.6 58.5 66.1 64.8
r/EDAnony 29.1 72.1 65.1 69.8 61.6 72.1 62.7 66.3
avg. diet 27.6 69.2 65.8 67.8 59.1 66.2 64.9 65.7
Table 1: Evaluation results of Llama-3.1-8B and GPT-3.5-Turbo on CommSurvey-test. The community-aligned LLMs (CrossInst and Steering+CrossInst) are finetuned on CommInst-train, so there is potential topic overlap between training and evaluation. For each model family, the results are divided into two groups, one without and one with steering. The best results in each group are highlighted in bold.
@@ 5.2 Baselines
LLM+Context.
For each community $C_{i}$, we provide the vanilla (unfinetuned) LLM with the context about the community by appending 300 most relevant comments from it. For example, when prompting the LLM to answer a question in CommSurvey-testi, we retrieve the most relevant 300 comments to the question, from the text chunks that are used to generate CommSurvey-traini, by calculating the embedding similarity with sentence-transformers. The prompt is augmented with the instruction “According to the following statements, learn the mindset and select only one most relevant answer by stating either A, B, C, or D. Do not provide any additional explanation or rationale for your choice. [comments]”. This baseline is inspired by the idea of in-context learning, where the LLM learns the community’s mindset within the context provided in the prompt.
Providing the LLM with context becomes less efficient when the model is deployed to answer a large number of survey questions, as each question requires processing an extremely long input sequence due to the added comments. This significantly increases computational costs, memory usage, and processing time, making it inefficient for large-scale or real-time applications. In contrast, our framework, which finetunes the model on community-specific data, is a one-time effort. Once the model is finetuned, it can efficiently answer survey questions without requiring additional long contextual inputs, making it far more scalable and resource-friendly in practical use cases.
LLM+Steering.
When prompting the vanilla LLM to answer survey questions, we steer it to mimic the community, by specifying in the prompt that “Select only one answer that best aligns with the opinions of members from subreddit r/[subreddit]”. Steering the vanilla LLMs can nudge them to respond to the community. For fair comparison to this baseline, we also steer the finetuned LLM via Community-Cross-Instruct, where the LLM is aligned to the community both in the finetuning and steering process.
It is worth noting that steering only applies to predefined communities that are developed via a manually specified tag, such as subreddits, where we can easily reference community names in the prompt. While this paper focuses on forum-based communities, Community-Cross-Instruct generalizes to other online communities, including organically formed communities in the retweet network Chu et al. (2024) or the news co-sharing network Mosleh et al. (2021); He et al. (2024b), as long as their relevant text data is readily available. Although it is not always feasible to concisely summarize organically formed communities in text, Community-Cross-Instruct allows for LLM alignment to such communities without requiring explicit textual descriptions.
Subreddit Llama-3.1-8B GPT-3.5-Turbo
Without Steering With Steering Without Steering With Steering
Context CrossInst Steering Steering+CrossInst Context CrossInst Steering Steering+CrossInst
r/Liberal 11.1 50.0 58.3 61.1 41.7 61.1 63.9 65.1
r/NeutralPol 20.8 60.0 54.2 65.0 52.5 65.0 67.5 67.5
r/Anar_Cap 23.9 56.5 52.2 56.5 54.3 58.7 58.7 71.7
r/Conservative 20.0 54.3 52.5 65.2 50.0 67.4 69.6 65.2
r/AT_Donald 25.0 56.5 57.5 56.5 52.2 60.9 61.2 63.0
avg. politics 20.2 55.5 54.9 60.9 50.1 62.6 64.2 66.5
r/keto 29.7 64.4 65.2 63.6 59.3 67.8 61.0 63.5
r/WLAdvice 26.9 63.5 61.5 62.5 58.7 62.5 63.5 62.5
r/EDAnony 25.4 55.9 50.0 54.9 57.8 60.8 58.8 61.8
avg. diet 27.3 61.3 58.9 60.3 58.6 63.7 61.1 62.6
Table 2: Evaluation results of Llama-3.1-8B and GPT-3.5-Turbo on CommSurvey-test-topic. The community-aligned LLMs (CrossInst and Steering+CrossInst) are finetuned on CommInst-train-topic, so there is no topic overlap between training and evaluation. For each model family, the results are divided into two groups, one without steering and one with steering. The best results in each group are highlighted in bold.
@@ 5.3 Main Results
We compare the finetuned LLMs’ generated answers to multi-choice questions to the “semi-ground truths” and report the accuracy in Figure 1 as a measure of alignment with the corresponding community. The LLMs are finetuned on CommInst-train and evaluated on CommInst-test.
Politics.
In the politics domain, CrossInst consistently outperforms Context across both LLMs, demonstrating the strength of finetuning on community-specific instructional data. By aligning LLMs to explicit community instructions, CrossInst enables the models to better capture the nuances of political discourse, where ideological divides and subtle differences in language are critical. This makes the model more adept at accurately reflecting community values without being overwhelmed by noisy or redundant information, which is a common issue with Context. We observe that the a large portion of generated answers from Llama-3.1 using Context are texts irrelevant to the survey questions, indicating that it struggles in dealing with a lengthy prompt with 300 examples.
An important aspect to consider is that LLMs come with pre-existing knowledge about various communities, which is learned during pretraining on large-scale internet data. During Steering, the model attempts to retrieve and apply this pre-existing knowledge to align its responses with the given community. However, this knowledge may not always be fully accurate or consistent with the actual values of the community being evaluated. For example, Steering might prompt the model to draw on generalizations or stereotypes that are present in its pretraining data, which might not reflect the current or specific views of the subreddit in question. This inconsistency can explain why Steering alone sometimes leads to suboptimal performance.
When combining Steering with CrossInst, a potential conflict arises between the knowledge gained from finetuning and the pre-existing community knowledge the model tries to retrieve through steering. The model is essentially being pulled in two directions: one based on the explicit, finetuned instructions from CrossInst and the other based on its internalized, sometimes outdated, pretraining knowledge. This conflict can result in Steering+CrossInst underperforming compared to Steering alone, as observed in certain subreddits, such as r/Anarcho_Capitalism on Llama-3.1. In these cases, the inconsistency between the steering prompts and the finetuned knowledge creates confusion for the model, leading to lower performance.
Diet.
In the diet domain, we observe a similar pattern where CrossInst significantly outperforms the Context baseline. Communities like r/keto and r/EDAnonymous, which are focused on specific health and lifestyle goals, benefit greatly from instructional finetuning. These communities are characterized by practical and focused discussions, and CrossInst allows the model to adapt to the specific language and norms within these subreddits, ensuring a more accurate reflection of their values and objectives.
Interestingly, in r/WeightLossAdvice, Steering alone performs the best across the four methods within each LLM. One possible reason for this is that in certain communities, particularly those that are well-represented in pretraining data, the model’s pre-existing knowledge might be more aligned with the actual community values. In such communities, where strong ideological markers may already be embedded in the model’s pretraining data, steering allows the model to retrieve this relevant information effectively, leading to stronger performance. In such cases, Steering helps amplify the model’s pre-learned alignment with the community, which can be sufficient for capturing the community’s voice without the need for additional finetuning.
@@ 5.4 Out-of-Topic Generalizability
Our final goal is to create LLMs aligned to different communities, which can be used to answer any survey question from the perspective of the community. In the real world, the surveys may contain questions covering topics that do not appear in the finetuning data CommInst. In this study, we finetune the LLMs on CommInst-train-topic, and evaluate them on CommSurvey-test-topic, to make sure that there is no overlap in topic coverage.
The results are shown in Table 2. CrossInst models continue to outperform the Context baselines, demonstrating their strong generalization capability to new topics, suggesting that CrossInst maintains its robustness even when exposed to entirely new topics.
We observe that Steering alone occasionally performs the best (e.g., r/AskThe_Donald on both LLMs, and r/WeightLossAdvice on GPT-3.5). We argue that this might be because ideologically-polarized or practical communities are consistent in their mindsets over time, so the LLM’s pre-knowledge about them would well predict their ideologies in the future.
@@ 6 Related Works
A survey of instruction-tuning approaches Zhang et al. (2023) outlines several methods for models to autonomously self-improve their instruction set. These include generating instructions for pre-existing texts Li et al. (2023), eliciting interaction between different model iterations Chen et al. (2024c), and bootstrapping from an existing instruction set to generate new ones Wang et al. (2022); Chen et al. (2024b). Additionally, a human-in-the-loop framework builds upon self-generalization by iterating between human and machine-generated instructions Guo et al. (2024). Despite these advances, these strategies require manual input. In this work, we build on self-instruct, a method to enhance instruction tuning by eliciting synthetic model instruction generations. We alter the original framework by removing the required set of manually written seed instructions Wang et al. (2022). This work significantly contributes to the field by offering a scalable and fully autonomous solution to instruction tuning, paving the way for more adaptive and intelligent models.
6.2 Aligning LLMs to Subgroups
Existing work has aligned LLMs to different human subgroups to discover their mindset and opinions Dorn et al. (2023, 2024). Subpopulation representative models (SRMs) Simmons and Hare (2023) can be used to emulate some characteristics of a particular subpopulation, particularly as LLMs can provide fine-grained, demographically-correlated outputs Argyle et al. (2023a). Argyle et al. (2023b) find that exposing GPT-3 to thousands of socio-demographic backstories leads the model to obtain a complex understandings of sociological concepts.
To learn about partisan communities, Jiang et al. (2022) propose CommunityLM by finetuning GPT-2 on tweets authored by prominent community members, and prompt the finetuned model to generate opinions about various political entities. He et al. (2024b) extends CommunityLM to organically-formed online communities with more fine-grained ideologies. However, these finetuned GPT-2 models can only be used for text continuation tasks and cannot answer survey questions. In our paper, we finetune the foundational instruction-tuned LLMs to serve as their digital twins. The resulting LLMs retain the instruction-following capabilities and are able to complete various tasks as specified in the instructions, including survey completion.
6.3 Evaluating LLMs’ Subgroup Alignment
Social scientists use surveys to systematically collect data from populations to characterize their beliefs, attitudes, opinions, and behaviors Hill et al. (1997); Choi and Pak (2005); Gordon (1987); Goel and Salganik (2010). LLM developers and practitioners focus on measuring LLM’s alignment with different human subgroups Santurkar et al. (2023); Durmus et al. (2023); He et al. (2024a) using use real-world survey responses as ground truth for comparing values. However, they often ignore under-represented groups, as it is difficult to administer surveys to people from those groups. To administer surveys, researchers sample individuals from a target population and ask them to manually respond to survey questions. This can be costly and slow, which limits the utility and timeliness of surveys. Our method addresses this challenge by using social media data that is easily accessible, even for those hard-to-reach subgroups. By automatically creating multi-choice survey questions and answers CommSurvey for these groups, our method enables a more comprehensive and cost-effective evaluation of LLMs’ alignment with diverse groups.
7 Discussions
We aim to align LLMs to online communities through instruction-tuning so that the aligned LLMs can be flexibly surveyed to probe the communities’ views on different topics. To prepare data for instruction tuning, we propose Community-Cross-Instruct, and the automatic creation of the open-ended instruction-response pairs (CommInst) is the core contribution. However, in order to demonstrate that the finetuned LLMs are indeed aligned to communities and can be treated as their digital twins, we propose to automatically generate survey questions and answers (CommSurvey) to efficiently evaluate LLM alignment. This is a secondary contribution of this paper, and we do not claim that CommSurvey is the best way to measure alignment. Therefore, our main goal is to propose a method to prepare LLMs as a proxy for surveying different online communities, rather than automatically creating such surveys.